OpenStudy (anonymous):

Suppose that F(x)=f(g(x)) and g(14)=2, g'(14)=5, f'(14)=15, and f'(2)=12. Find F'(14.)

OpenStudy (anonymous):

F'(x) = f'(g(x)) * g'(x) so F'(14) = f'(g(14)) * g'(14) now substitute all those givens to get your answer...

OpenStudy (anonymous):

umm.. not quite sure what you mean.

OpenStudy (anonymous):

15(2) * 5

OpenStudy (anonymous):

it's the chain rule.... F'(14) = f'(g(14)) * g'(14) = f'(2) * 5 = 12 * 5 = 60

OpenStudy (anonymous):

oh cool! i get it..

OpenStudy (anonymous):

got a few more to do for this one

OpenStudy (anonymous):

yupp! haha

OpenStudy (anonymous):

ok... close shop again... post em...